Please help.
Solve 2/5x=-10

Answer by checkley71(8403)

IF YOU MEAN (2/5)X=-10 THEN:
2X/5=-10 NOW CROSS MULTIPLY
2X=-50
X=-50/2
X=-25 ANSWER.
PROOF
(2/5)-25=-10
-50/5=-10
-10=-10